James Tetazoo

Term Address Home Address

3 Ames Street 3 Totters Field Rd. Cambridge, MA 02139 Freeport, ME 04033 (617) 225-6378

Employment A position involving computer systems security Objective

Education Massachusetts Institute of Technology 1986-present Candidate for the S.B. degree in Computer Science and Engineering, June 1990. Thesis work will be on “Provably Insecure Systems.” Relevant courses taken include Automata, Computability, and Complexity, Intorduction to Algorithms, Software Engineering Lab, Telephony, Computer Systems Engineering, and Cryptography and cryptanalysis. GPA: 4.9/5.0

Honors Membership in the following: Mensa, Titan, and the Triple Nine Society.

Member of Eta Kappa Nu and Tau Beta Pi.

Experience MIT Information Services Cambridge, MA UROP Student (August 1988-May 1989)—worked with Professor Eichin evaluating security of popular network software.

Charles Inman Anderson Co., Ltd. Cambridge, MA Summer Student (May-August 1988, June-August 1989)—evaluation and testing of real world security systems, both electronic and physical.

MIT Department of Architecture Cambridge, MA UROP Student (August 1987-May 1988)—development programming and testing for a computer software package to track building maintenance for physical plant.

MIT Dormline Cambridge, MA Student Worker (August 1986-May 1987)—repair and maintenance of central switching facilities for MIT’s dormitory phone system.

Joe’s Locks and Hardware Freeport, ME Apprentice Locksmith (May-August 1986, June-August 1987)—routine repair and rekeying of all popular locks. Occasional field work for customers who had lost keys.

Knowledge of Pascal, C, Scheme, and 680x0 an VAX Assembler. Familiarity with Unix (Berkeley and System V), TCP/IP, Kerberos, and NFS.

Spelunker’s Club Cambridge, MA Activities and Interests

Fall 1986-present—led and helped with urban spelunking within the MIT community.

THiRD EAst Travelling Animal Zoo Cambridge, MA Fall 1986-present—helped plan and participate in numerous jokes, water fights, and hacks. Recent involvement includes instrumental roles in the construction and creative deployment of bats, spaceships, and stop lights, as well as the semi-annual soaking of Larry.